OpenStudy (anonymous):

Please can any one help? According to genetic theory, plants of a particular species have a 25% chance of being red-flowering, independently of other plants. Find the normal approximation to the chance that among 10,000 plants of this species, more than 2400 are red-flowering. Approach:You want P(X>2400) when X follows a binomial distribution with p=0.25 and N=10000. You need a normal approximation Y to X - that is, you should match the mean and variance of Y to those of X. Then you can approximate your answer using P(Y>2400)?

OpenStudy (agent0smith):

Find P(X>2400), using P(Y>2400) as a normal approximation. First find the mean for the population of 10000. Since p=0.25, the expected amount (mean) of red flowering plants is: \[\Large E(x) = np = 0.25*10000 = 2500\] Standard deviation for the normal approximation is given by \[\Large SD = \sqrt{ np (1-p)} = \sqrt{ 2500 (1-0.25)} = 43.301\] Now we can find the z-score so we can find P(Y>2400). z score is given by \[\Large z = \frac{ X - mean }{ SD }\] \[\Large z = \frac{ 2400 - 2500 }{ 43.301 } = -2.31\] Now you just have to use a z chart to solve: http://www.mathsisfun.com/data/standard-normal-distribution-table.html You need P(z>-2.31)...ie the area to the right of -2.31

OpenStudy (agent0smith):

Or: http://www.wolframalpha.com/input/?i=P%28z%3E-2.31%29

OpenStudy (agent0smith):

If you're confused, this site has step by step directions to a similar problem: http://www.statisticshowto.com/articles/using-the-normal-approximation-to-solve-a-binomial-problem/ I didn't use the continuity correction factor, but it makes ~no difference in this question anyway.

OpenStudy (anonymous):

Answer is 0.9896 - (if somebody has problem with "SticiGui")

OpenStudy (anonymous):

is the answer is 98.96% or 0.9896

OpenStudy (agent0smith):

One of them is a percentage probability, the other is a decimal probability...

OpenStudy (anonymous):

the lower limit -2.31 and upper limit 5 area under the normal curve =98.96%=0.9896 is slot answer

OpenStudy (anonymous):

thanks agent0smith

OpenStudy (anonymous):

0.9896 - decimal is right (without %)

OpenStudy (anonymous):

Thanks Oles if You enter in percent 98.96 You get it wrong so its correct to enter always in slot answer 0.9896!

OpenStudy (agent0smith):

@Oles 98.96% is also correct. Just depends on what answer the test wants, a percentage or decimal. They're equivalent.

OpenStudy (anonymous):

in our case for this type it wants decimal agent0smith thanks both are correct!

OpenStudy (anonymous):

Hi! I will put to my attantion that answer is accept both formats: percentage and decimal. In any case its extra experiance for me ;-) Thanks